In a college, 4/9 of all employees are faculty and the ratio of the faculty to support staff is 5:4. How many people does the college employ if the support staff numbers 192?

Answer by rfer(16322)

192/4*5=240 faculty
240/(4/9)=
240*9/4=540 total